OpenStudy (anonymous):

squared x+8=x-4 a. 8 b. -1 c. -4,4 d. no real number solution

OpenStudy (anonymous):

8 -4 so 4?

OpenStudy (anonymous):

c is the answer?

OpenStudy (anonymous):

D ithink

OpenStudy (anonymous):

why?

OpenStudy (jhannybean):

?... \(x^2 -8 = x+4\) is that your question?..

OpenStudy (anonymous):

√x+8=x-4

OpenStudy (jhannybean):

Oh, \[\sqrt{x-8}=x+4\] Start by squaring both sides \[(\sqrt{x-8})^2 = (x+4)^2\]

OpenStudy (anonymous):

A

OpenStudy (anonymous):

X+8=X^2+16-8X

OpenStudy (anonymous):

X^2-9X+8=0

OpenStudy (anonymous):

DELTA=81-32=49

OpenStudy (anonymous):

so it would be a?

OpenStudy (anonymous):

X1++9+7/2=8

OpenStudy (jhannybean):

When you square both sides, the square root goes away,leaving you with:\[(\sqrt{x-8})^2 = (x+4)^2\]\[x-8 = x^2 +8x+16\]

OpenStudy (anonymous):

,X2=9-7/2=1

OpenStudy (anonymous):

IT'S A

OpenStudy (anonymous):

@Jhannybean (X-4) NOT (X+4)

OpenStudy (anonymous):

thanks guys can u help me with more?

OpenStudy (jhannybean):

Oop, typo.

OpenStudy (anonymous):

YEP

OpenStudy (anonymous):

should i open a new question?

OpenStudy (anonymous):

YEP